OpenStudy (anonymous):

How do I Prove or disprove; For all x, y € R, if x is rational and y is irrational, then xy is irrational

OpenStudy (2bornot2b):

\(x\) is rational, so \(x\) can be expressed as \(\frac{p}{q}\) Let us assume that \(x\times y\) is rational, then \(xy=\frac{m}{n}\implies y=\frac{mq}{np}\) . This shows that y is rational, which is a contradiction. So \(xy\) is irrational.

OpenStudy (2bornot2b):

Did you understand it?